OpenStudy (anonymous):

20. 70 = –7(–2 – 2z) (1 point)

OpenStudy (anonymous):

70 = -7(-2-2z) 70 = 14 +14z 56 = 14z z = 56/14

OpenStudy (anonymous):

Ask if you're unsure why.

OpenStudy (anonymous):

That makes since, thanks!!

OpenStudy (anonymous):

sense! sorry lol